OpenStudy (anonymous):

Find dy/dx of y=sin^(-1)x + cos^(-1)x

OpenStudy (anonymous):

\[y=\sin ^{-1}x+\cos^{-1}x\]

hartnn (hartnn):

use the identity that sin^(-1)x + cos^(-1)x is just pi/2 and dy/dx of y=pi/2 is ?

OpenStudy (anonymous):

i got \[\frac{ 1 }{ 2\sqrt{1-x^2} }\]

OpenStudy (anonymous):

But its the wrong answer in the book.

OpenStudy (anonymous):

wondering what i did wrong.

hartnn (hartnn):

\(\large \arcsin x +\arccos x =\pi/2\)

OpenStudy (anonymous):

How?

hartnn (hartnn):

its an identity http://en.wikipedia.org/wiki/Inverse_trigonometric_functions

OpenStudy (anonymous):

so you wont use the derivative formula?

hartnn (hartnn):

if you use derivative formula, you would still get a 0

OpenStudy (anonymous):

because i hate memorizing the identities, cant you use \[\sin^{-1}x=\frac{ 1 }{ \sqrt{1-x^2} }\]

hartnn (hartnn):

1/ sqrt (1-x^2) + (-1/ sqrt (1-x^2)) = 0

OpenStudy (anonymous):

Then can u teach me the long way by using the derivative formula?

OpenStudy (anonymous):

OHHHHHh ommmggggg that negative in cosine, totally didnt see that. ICIC thanks

hartnn (hartnn):

welcome ^_^